OpenStudy (anonymous):

d/dt (1-tan t/sec t)

OpenStudy (anonymous):

you can re-write as \[1- \frac{ \sin(t) \cos(t) }{ \cos(t) }\]

OpenStudy (anonymous):

aka the derivative of -sin(t)

OpenStudy (anonymous):

what about 1/(1/cos)?

OpenStudy (anonymous):

=cos

OpenStudy (anonymous):

shouldn't is be cos-sin?

OpenStudy (anonymous):

*it

OpenStudy (anonymous):

why? because the problem is actually \[\frac{ d}{dt }\frac{ 1-\tan t }{\sec t }\] ?

OpenStudy (anonymous):

i get it now, sorry about that! thanks again